# TransformerFAM: Feedback attention is working memory

### v2

@@ 1 Introduction
(a) Sliding Window Attention
@@
Figure 2: Comparison of attention patterns in Transformer layer. (a) TransformerBSWA: Input query attends to the current block and two memory segments, providing past context. (b) TransformerFAM: Input query attends to the current block, memory segments, and past FAM (green lines). FAM query (copied from previous FAM, blue dash arrow) compresses the current block to update FAM. This feedback loop enables information compression and propagation over indefinite horizon, which is working memory. Figure 5 shows in detail how the dynamic process occurs over time.
The introduction of the Transformer architecture (Vaswani et al., 2017) has revolutionized deep learning by permeating diverse domains and enhancing performance due to its efficacy and scalability. This scalability fuels a trend analogous to Moore’s law, which links increased model size to performance gains (Kaplan et al., 2020).
The effectiveness of attention in text sequence processing was solidified through the Transformer paper. Models like BERT (Devlin et al., 2018) and GPT-3 (Brown et al., 2020) further showcased the scalability of Transformer and its tendency for improved performance with increased model size. Following the replacement of LSTM (Hochreiter & Schmidhuber, 1997b) by Transformer in most Natural Language Processing (NLP) domains, the Vision Transformer (ViT) (Dosovitskiy et al., 2020) replaced Convolutional Neural Network (CNN) (LeCun et al., 1995) with Transformers in the vision domain, and Conformer (Convolution-augmented Transformer) (Gulati et al., 2020) replaced LSTM in the speech domain. The Transformer has become the de facto architecture in various domains. Currently, attention serves as the leading architecture for extracting meaningful representations from homogeneous data.
The logical progression points toward extending attention capabilities to heterogeneous data. This has enabled advances in multimodal fusion (text and vision), as seen in models like DALL·E 2 (Ramesh et al., 2022), Flamingo (Alayrac et al., 2022) and CoCa (Yu et al., 2022). AudioLM (Borsos et al., 2023) has shown that attention also excels at fusing audio and text. Consequently, Gemini (Team et al., 2023) integrates text, images, audio, and video into a single generative model. This was possible because attention to heterogeneous data works exceptionally well.
@@
Sliding window attention is introduced (Dai et al., 2019; Beltagy et al., 2020) to handle infinitely long sequences as input. However, it disregards information beyond the effective receptive field (approximately model depth $\times$ window size in Figure 1(d)) during sequence generation. Various approaches have attempted to address the long-context problem by sparse attention (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020; Kitaev et al., 2020; Roy et al., 2021; Oren et al., 2024) and linear approximated attention (Wang et al., 2020; Katharopoulos et al., 2020; Choromanski et al., 2020; Xiong et al., 2021), showing effectiveness below the 1B scale. Yet, scaling laws (Tay et al., 2022) suggest these approximations do not perform as well at GPT-3 level. Current SoTA large language model (LLM) leaderboards (LMSYS, 2023) do not feature architectures primarily relying on approximated attention.
On the other hand, Neuroscience links attention to multisensory integration (Tang et al., 2016). Endogenous (goal-driven) and exogenous (stimulus-driven) attention are distributed throughout sensory processing regions in brain, and the brain areas correlated with the attention overlap substantially with the multisensory cortical regions involved in multisensory integration. Working memory exhibits similar distribution (Christophel et al., 2017), overlapping with attention-correlated areas. This indicates that attention is an important factor in not only multimodal fusion but also working memory.
In the human brain, working memory (Wallace, 1960) provides a temporary memory for performing tasks. While working memory is stored in sustained activations, long-term memory is stored in weights (Fuster, 1973). LLMs has enormous long-term memory enough to store the entire internet (Villalobos et al., 2022), but they do not have working memory. The activation of working memory is sustained by prefrontal cortical-thalamic loops (Ashby et al., 2005), which means working memory is sustained by the continuous spiking of activation within a feedback loop.
@@
Feedback connections are prevalent in biological neural networks. Even organisms with simple neural structures, such as C. elegans (with only 302 neurons) (White et al., 1986), exhibit various feedback loops, like connections from higher-level interneurons to lower-level ones (Hasani et al., 2018). However, incorporating feedback loops in Transformer is challenging. There are two main approaches. The first approach is linking the topmost layer to the bottommost (Fan et al., 2020; Bulatov et al., 2022). However, this cannot model feedback between interneurons, and this has only one global working memory. The second approach is within-transformer-block feedback: the output activation of a Transformer layer is fed back as input to the same layer. This is the approach we propose, which enables each Transformer layer to have a distributed working memory that corresponds to its abstraction level.
Recurrent Neural Networks (RNNs) have achieved great success in machine learning by introducing feedback loops (Hochreiter & Schmidhuber, 1997a; Cho et al., 2014). RNNs pass feedback between sequences through hidden states. Attention mechanisms can implement feedback loops by attending to both the input sequence and the feedback state simultaneously.
@@
TransformerFAM does not introduce new weights to the Transformer, allowing the reuse of pre-trained checkpoints. Our experiments show that fine-tuning TransformerFAM with LoRA for just 50k steps significantly enhances performance on long-context tasks across 1B, 8B, and 24B Flan-PaLM LLMs (Chung et al., 2022).
@@ 2.1 Block Sliding Window Attention (BSWA)
In general, when handling long context inputs, there are the two main approaches. A first approach is to increase the context length with increasing the computational resources (memory and processing power). A second approach is the implementation of Sliding Window Attention (SWA) (Beltagy et al., 2020; Jiang et al., 2023), as illustrated in Figure 1(a). During inference, the standard implementation allocates key and value cache twice the window length at the beginning, and then using a ring buffer to update the necessary components at each step, in order to avoid memory allocation and copying operations every step, which are computationally expensive.
Longformer (Beltagy et al., 2020) introduced Sliding Window Attention, which caches on a block-by-block basis. We will refer to this as Block Sliding Window Attention (BSWA). BSWA does not mask out past keys and values in the ring buffer, but attends to all the information in the ring buffer.
@@
Given a sequence input $I=[i_{1},i_{2},\dots,i_{T}]$ to a vanilla Transformer layer, Transformer transforms the input to a sequence output $O=[o_{1},o_{2},\dots,o_{T}]$ as follows (Vaswani et al., 2017):
@@
While PreLN (Xiong et al., 2020) is used in the equation, it is not mandatory. Autoregressive Transformer changes the SelfAttention equation as follows:
$a_{t}$ $=\text{SelfAttention}(q_{t},K_{:t},V_{:t})+i_{t}$
In a Transformer with SWA, the equation is modified to account for a window size of $w$:
$a_{t}$ $=\text{SelfAttention}(q_{t},K_{t-w:t},V_{t-w:t})+i_{t}$
In a Transformer with BSWA, $\tau$ denotes the block index. Each block $\tau$ contains a set of keys and values, determined by the block size. The equation is modified to account for a block index $\tau$ and a memory segment $m$ in Equation 1. $K_{\tau-m:\tau-1}$ is the keys in the memory segments from the $m$ blocks before $\tau$ to the block before $\tau$, and $K_{\tau,:t}$ is the keys from the beginning of $\tau$ block up to $t$. We will refer to Transformer with BSWA as TransformerBSWA.
@@
Algorithm 1 presents TransformerBSWA (Equation 1), re-expressed from the perspective of a block index $\tau$. Algorithm 1 describes how to iteratively calculate $a_{t}$ and then concatenate the results into an $A_{\tau}$ sequence. Typically, standard implementations employ a causal attention mask to enable parallel computation of self-attention.
TransformerXL (Dai et al., 2019) proposed to use a technique called ”stop gradient” for the memory segment. However, we argue that this technique has a negative impact on ability of the model to attend to past information. Specifically, we show that using stop gradient results in a much shorter receptive field than the theoretical receptive field in Section 3.2. We believe that allowing gradient to flow to the memory segment is necessary for the model to learn to carry important information in the memory segment.
@@
While this modification might seem to burden training memory and computation, it does not significantly impact performance in practice. This is primarily due to the prevalence of gradient checkpointing (Chen et al., 2016) in LLM training on ML accelerators, as memory often presents the primary bottleneck. Gradient checkpointing recomputes attention during backpropagation, from later blocks to earlier blocks. Therefore, the presence or absence of stop gradients has little impact on the overall computational complexity, while still improving performance.
@@
Transformer has quadratic memory and computation complexity with respect to the length of the input sequence due to self-attention. It has $O(L^{2})$ complexity for input length $L$. However, Transformer with Sliding Window Attention has linear complexity with respect to the input sequence. It has $O(L\times W)$ complexity for input length $L$ and window size $W$. If the input length is about the same as the window size ($\sim 1$k), the complexity difference is almost negligible, but if the input length is large like $128$k in GPT-4 turbo (Achiam et al., 2023), there is a huge difference. In inference, Transformer with SWA or BSWA only needs to cache a fixed ring buffer (block size $+$ memory segment). Therefore, it only consumes constant memory regardless of the generated token length. Therefore, LLMs using SWA or BSWA can generate infinitely long output tokens.
However, BSWA has a limited receptive field, approximately equal to model depth $\times$ window size as illustrated in Figure 1(d). As a result, the later generated tokens are not related to tokens outside the receptive field (e.g., prompt). To address this limitation, we propose a novel architecture in the following Section 2.2. Our approach, Feedback Attention Memory (FAM), builds upon BSWA. This is because the block stride concept of BSWA is well-suited for blockwise feedback updates.
@@ 2.2 Feedback Attention Memory
As mentioned in Section 1, we hypothesized that attending to the feedback loop can give rise to working memory in 1.1. To implement the feedback loop, we add feedback activations that feed contextual representation back into each block of BSWA. We call these virtual activations as Feedback Attention Memory (FAM). FAM is designed to meet the following key requirements:
@@
The proposed architecture achieves this by appending FAM to block segments and incorporating it into self-attention processes. This enables richer representations and dynamic propagation of global contextual information across blocks, as illustrated in Figure 2(b). When self-attention occurs on the current block, the input query for the block attends to the input key for that block, the memory segment, and the previous FAM. The previous FAM provides global contextual information, allowing for a much richer representation than BSWA. In parallel, the FAM query attends to the current block and the FAM key. The FAM query compresses the current block, conditioned on the previous global contextual information. The FAM query is dynamically generated based on previous global contextual information, as it is copied from the previous FAM. Then, the newly updated FAM serves to propagate global contextual information to the next block recursively. This process is formally described in Algorithm 2.
@@ 3.1 Training
Pretraining an LLM from scratch requires a huge amount of resources. TransformerFAM can reuse existing LLM checkpoints because it does not add new weights to the Transformer layer. We reused 1B, 8B, and 24B Flan-PaLM LLMs (Chung et al., 2022) for our experiments. This is a large enough size to prove that TransformerFAM is a general solution for LLMs. The model sizes 1B, 8B, and 24B refer to the size of the plain Transformer, excluding the text embedding table. The models use a 256k sentence piece tokenizer (Kudo & Richardson, 2018), resulting in 400M text embedding table weights for the 1B model and 1B weights for the 8B and 24B models. The detailed model architecture is described in Table 5 in Section A.1.
Flan-PaLM is a model that is fine-tuned on top of a pretrained PaLM model (Chowdhery et al., 2023) using instruction finetuning. The instruction data consists of few-shot instructions with $100$ to $1$k tokens, which are packed into $2.5$k tokens for training. This means that individual instruction data are concatenated until they reach $2.5$k tokens.
@@
During fine-tuning, we used the same Flan instruction data packed into $8.5$k tokens. To maintain a minibatch size of 128 for all models, we used 32 TPUv5 (Jouppi et al., 2023) cores for the 1B model, 64 cores for the 8B model, and 128 cores for the 24B model. If we had used more resources and a larger minibatch size, we might have achieved better results than those reported in the paper.
We performed LoRA finetuning by adding LoRA (Hu et al., 2021) to the Attention and FF layers of Transformer without training all the parameters. Full finetuning resulted in lower scores on various tasks reported by GPT-3 (Brown et al., 2020), because catastrophic forgetting (Kirkpatrick et al., 2017) occurred in domains that were not covered by the instruction data. In LoRA finetuning, the scores on GPT-3 tasks actually improved, and the performance on long context tasks was similar to that of full finetuning. The rank of LoRA was 64, and the weights of the original Attention and FF layers were merged with LoRA weights and used during inference.
The Adafactor optimizer ( $\beta_{1}=0.9,\beta_{2}=0.99$ )  (Shazeer & Stern, 2018) was used with constant learning rate. The learning rates used were $10^{-4}$ for 1B, and $3\times 10^{-3}$ for both 8B and 24B.
In addition, TransformerXL exhibits comparable performance to TransformerBSWA. The implementations are almost identical, with TransformerXL employing an additional QK attention mask to mask out keys beyond a predetermined window size. Section B.4.5 demonstrates that the performance difference between the two architectures is insignificant, and therefore, experimental results for TransformerBSWA are only included in the main paper.
@@ 3.1.1 Data
We used Flan instruction data as training data, packed up to 8.5k tokens. In the Flan paper (Chung et al., 2022), a special attention mask was used to prevent attention between different segments during self-attention, by applying a separate mask to each packed segment. We did not use this special attention mask processing. Attention occurs causally within the window, regardless of the segment. We expected TransformerBSWA and TransformerFAM to learn to remember important information and forget outdated information by themselves.

### v3

@@ 1 Introduction
The introduction of the Transformer architecture [12] has revolutionized deep learning by permeating diverse domains and enhancing performance due to its efficacy and scalability. This scalability fuels a trend analogous to Moore’s law, which links increased model size to performance gains [39].
The effectiveness of attention in text sequence processing was solidified through the Transformer paper. Models like BERT [16] and GPT-3 [33] further showcased the scalability of Transformer and its tendency for improved performance with increased model size. Following the replacement of LSTM [5] by Transformer in most Natural Language Processing (NLP) domains, the Vision Transformer (ViT) [32] replaced Convolutional Neural Network (CNN) [4] with Transformers in the vision domain, and Conformer (Convolution-augmented Transformer) [29] replaced LSTM in the speech domain. The Transformer has become the de facto architecture in various domains. Currently, attention serves as the leading architecture for extracting meaningful representations from homogeneous data.
The logical progression points toward extending attention capabilities to heterogeneous data. This has enabled advances in multimodal fusion (text and vision), as seen in models like DALL·E 2 [55], Flamingo [54] and CoCa [53]. AudioLM [64] has shown that attention also excels at fusing audio and text. Consequently, Gemini [69] integrates text, images, audio, and video into a single generative model. This was possible because attention to heterogeneous data works exceptionally well.
(a) SWA
@@
Figure 2: Comparison of attention patterns in Transformer layer. (a) TransformerBSWA: Input query attends to the current block and two memory segments, providing past context. (b) TransformerFAM: Input query attends to the current block, memory segments, and past FAM (green lines). FAM query (copied from previous FAM, blue dash arrow) compresses the current block to update FAM. This feedback loop enables information compression and propagation over indefinite horizon, which is working memory. Fig. 4 shows in detail how the dynamic process occurs over time.
@@
Sliding window attention is introduced [26, 35] to handle infinitely long sequences as input. However, it disregards information beyond the effective receptive field (approximately model depth $\times$ window size in Fig. 1(d)) during sequence generation. Various approaches have attempted to address the long-context problem by sparse attention [23, 35, 28, 37, 42, 81] and linear approximated attention [34, 40, 38, 48], showing effectiveness below the 1B scale. Yet, scaling laws [59] suggest these approximations do not perform as well at GPT-3 level. Current SoTA large language model (LLM) leaderboards [82] do not feature architectures primarily relying on approximated attention.
On the other hand, Neuroscience links attention to multisensory integration [10]. Endogenous (goal-driven) and exogenous (stimulus-driven) attention are distributed throughout sensory processing regions in brain, and the brain areas correlated with the attention overlap substantially with the multisensory cortical regions involved in multisensory integration. Working memory exhibits similar distribution [13], overlapping with attention-correlated areas. This indicates that attention is an important factor in not only multimodal fusion but also working memory.
In the human brain, working memory [83] provides a temporary memory for performing tasks. While working memory is stored in sustained activations, long-term memory is stored in weights [2]. LLMs has enormous long-term memory enough to store the entire internet [62], but they do not have working memory. The activation of working memory is sustained by prefrontal cortical-thalamic loops [7], which means working memory is sustained by the continuous spiking of activation within a feedback loop.
@@
Feedback connections are prevalent in biological neural networks. Even organisms with simple neural structures, such as C. elegans (with only 302 neurons) [3], exhibit various feedback loops, like connections from higher-level interneurons to lower-level ones [17]. However, incorporating feedback loops in Transformer is challenging. There are two main approaches. The first approach is linking the topmost layer to the bottommost [27, 57]. However, this cannot model feedback between interneurons, and this has only one global working memory. The second approach is within-transformer-block feedback: the output activation of a Transformer layer is fed back as input to the same layer. This is the approach we propose, which enables each Transformer layer to have a distributed working memory that corresponds to its abstraction level.
Recurrent Neural Networks (RNNs) have achieved great success in machine learning by introducing feedback loops [6, 9]. RNNs pass feedback between sequences through hidden states. Attention mechanisms can implement feedback loops by attending to both the input sequence and the feedback state simultaneously.
@@
TransformerFAM does not introduce new weights to the Transformer, allowing the reuse of pre-trained checkpoints. Our experiments show that fine-tuning TransformerFAM with LoRA for just 50k steps significantly enhances performance on long-context tasks across 1B, 8B, and 24B Flan-PaLM LLMs [58].
@@ 2.1 Block Sliding Window Attention (BSWA)
In general, when handling long context inputs, there are the two main approaches. A first approach is to increase the context length with increasing the computational resources (memory and processing power). A second approach is the implementation of Sliding Window Attention (SWA) [35, 74], as illustrated in Fig. 1(a). During inference, the standard implementation allocates key and value cache twice the window length at the beginning, and then using a ring buffer to update the necessary components at each step, in order to avoid memory allocation and copying operations every step, which are computationally expensive.
Longformer [35] introduced Sliding Window Attention, which caches on a block-by-block basis. We will refer to this as Block Sliding Window Attention (BSWA). BSWA does not mask out past keys and values in the ring buffer, but attends to all the information in the ring buffer.
@@
Given a sequence input $I=[i_{1},i_{2},\dots,i_{T}]$ to a vanilla Transformer layer, Transformer transforms the input to a sequence output $O=[o_{1},o_{2},\dots,o_{T}]$ as follows [12]:
@@
While PreLN [36] is used in the equation, it is not mandatory. Autoregressive Transformer changes the SelfAttention equation as follows: $a_{t}=\text{SelfAttention}(q_{t},K_{:t},V_{:t})+i_{t}$.
In a Transformer with SWA, the equation is modified to account for a window size of $w$: $a_{t}=\text{SelfAttention}(q_{t},K_{t-w:t},V_{t-w:t})+i_{t}$.
In a Transformer with BSWA, $\tau$ denotes the block index. Each block $\tau$ contains a set of keys and values, determined by the block size. The equation is modified to account for a block index $\tau$ and a memory segment $m$ in Eq. 1. $K_{\tau-m:\tau-1}$ is the keys in the memory segments from the $m$ blocks before $\tau$ to the block before $\tau$, and $K_{\tau,:t}$ is the keys from the beginning of $\tau$ block up to $t$. We will refer to Transformer with BSWA as TransformerBSWA.
@@
Algorithm 1 presents TransformerBSWA (Eq. 1), re-expressed from the perspective of a block index $\tau$. Algorithm 1 describes how to iteratively calculate $a_{t}$ and then concatenate the results into an $A_{\tau}$ sequence. Typically, standard implementations employ a causal attention mask to enable parallel computation of self-attention.
TransformerXL [26] proposed to use a technique called "stop gradient" for the memory segment. However, we argue that this technique has a negative impact on ability of the model to attend to past information. Specifically, we show that using stop gradient results in a much shorter receptive field than the theoretical receptive field in Section 3.2. We believe that allowing gradient to flow to the memory segment is necessary for the model to learn to carry important information in the memory segment.
@@
While this modification might seem to burden training memory and computation, it does not significantly impact performance in practice. This is primarily due to the prevalence of gradient checkpointing [11] in LLM training on ML accelerators, as memory often presents the primary bottleneck. Gradient checkpointing recomputes attention during backpropagation, from later blocks to earlier blocks. Therefore, the presence or absence of stop gradients has little impact on the overall computational complexity, while still improving performance.
@@
Transformer has quadratic memory and computation complexity with respect to the length of the input sequence due to self-attention. It has $O(L^{2})$ complexity for input length $L$. However, Transformer with Sliding Window Attention has linear complexity with respect to the input sequence. It has $O(L\times W)$ complexity for input length $L$ and window size $W$. If the input length is about the same as the window size ($\sim 1$k), the complexity difference is almost negligible, but if the input length is large like $128$k in GPT-4 turbo [70], there is a huge difference. In inference, Transformer with SWA or BSWA only needs to cache a fixed ring buffer (block size $+$ memory segment). Therefore, it only consumes constant memory regardless of the generated token length. Therefore, LLMs using SWA or BSWA can generate infinitely long output tokens.
However, BSWA has a limited receptive field, approximately equal to model depth $\times$ window size as illustrated in Fig. 1(d). As a result, the later generated tokens are not related to tokens outside the receptive field (e.g., prompt). To address this limitation, we propose a novel architecture in the following Section 2.2. Our approach, Feedback Attention Memory (FAM), builds upon BSWA. This is because the block stride concept of BSWA is well-suited for blockwise feedback updates.
@@ 2.2 Feedback Attention Memory
As mentioned in Section 1, we hypothesized that attending to the feedback loop can give rise to working memory in 1. To implement the feedback loop, we add feedback activations that feed contextual representation back into each block of BSWA. We call these virtual activations as Feedback Attention Memory (FAM). FAM is designed to meet the following key requirements:
@@
The proposed architecture achieves this by appending FAM to block segments and incorporating it into self-attention processes. This enables richer representations and dynamic propagation of global contextual information across blocks, as illustrated in Fig. 2(b). When self-attention occurs on the current block, the input query for the block attends to the input key for that block, the memory segment, and the previous FAM. The previous FAM provides global contextual information, allowing for a much richer representation than BSWA. In parallel, the FAM query attends to the current block and the FAM key. The FAM query compresses the current block, conditioned on the previous global contextual information. The FAM query is dynamically generated based on previous global contextual information, as it is copied from the previous FAM. Then, the newly updated FAM serves to propagate global contextual information to the next block recursively. This process is formally described in Algorithm 2.
@@ 3.1 Training
Pretraining an LLM from scratch requires a huge amount of resources. TransformerFAM can reuse existing LLM checkpoints because it does not add new weights to the Transformer layer. We reused 1B, 8B, and 24B Flan-PaLM LLMs [58] for our experiments. This is a large enough size to prove that TransformerFAM is a general solution for LLMs. The model sizes 1B, 8B, and 24B refer to the size of the plain Transformer, excluding the text embedding table. The models use a 256k sentence piece tokenizer [20], resulting in 400M text embedding table weights for the 1B model and 1B weights for the 8B and 24B models. The detailed model architecture is described in Table 3 in Section A.1.
Flan-PaLM is a model that is fine-tuned on top of a pretrained PaLM model [75] using instruction finetuning. The instruction data consists of few-shot instructions with $100$ to $1$k tokens, which are packed into $2.5$k tokens for training. This means that individual instruction data are concatenated until they reach $2.5$k tokens.
@@
During fine-tuning, we used the same Flan instruction data packed into $8.5$k tokens. To maintain a minibatch size of 128 for all models, we used 32 TPUv5 [77] cores for the 1B model, 64 cores for the 8B model, and 128 cores for the 24B model. If we had used more resources and a larger minibatch size, we might have achieved better results than those reported in the paper.
We performed LoRA finetuning by adding LoRA [46] to the Attention and FF layers of Transformer without training all the parameters. Full finetuning resulted in lower scores on various tasks reported by GPT-3 [33], because catastrophic forgetting [14] occurred in domains that were not covered by the instruction data. In LoRA finetuning, the scores on GPT-3 tasks actually improved, and the performance on long context tasks was similar to that of full finetuning. The rank of LoRA was 64, and the weights of the original Attention and FF layers were merged with LoRA weights and used during inference.
The Adafactor optimizer ( $\beta_{1}=0.9,\beta_{2}=0.99$ )  [15] was used with constant learning rate. The learning rates used were $10^{-4}$ for 1B, and $3\times 10^{-3}$ for both 8B and 24B.
In addition, TransformerXL exhibits comparable performance to TransformerBSWA. The implementations are almost identical, with TransformerXL employing an additional QK attention mask to mask out keys beyond a predetermined window size. Section B.4.7 demonstrates that the performance difference between TransformerXL and TransformerBSWA is insignificant, and therefore, experimental results for TransformerBSWA are only included in the main paper.
@@ 3.1.1 Data
We used Flan instruction data [84] as training data, packed up to 8.5k tokens. In the Flan paper [44, 58], a special attention mask was used to prevent attention between different segments during self-attention, by applying a separate mask to each packed segment. We did not use this special attention mask processing. Attention occurs causally within the window, regardless of the segment. We expected TransformerBSWA and TransformerFAM to learn to remember important information and forget outdated information by themselves.